OpenStudy (anonymous):

What is the 22nd term of the arithmetic sequence where a1 = 8 and a9 = 56?

OpenStudy (the_fizicx99):

The arithmetic formula is: a_n = a1 + d(n-1) First find the difference, 56 - 8 = 48/ 8 = 6 so d = 6 a_22 = 8 + 6(22 - 1) a_22 = 8 + 6(21) a_22 = 8 + 126 a_22 = 134

OpenStudy (anonymous):

Thank you so much!

OpenStudy (the_fizicx99):

Yw